?? 1 ∼ 1^{∘} 2048× 4096 ?? ∼ 3× ∼ 3 ??

R ∼ 40′′ ∼ 21 ∼ 15

Based on photographic data obtained using Oschin Schmidt Telescope on Palomar Mountain. The Palomar Observatory Sky Survey was funded by the National Geographic Society. The Oschin Schmidt Telescope is operated by the California Institute of Technology and Palomar Observatory. The plates were processed into the present compressed digital format with their permission. The Digitized Sky Survey was produced at the Space Telescope Science Institute (ST ScI) under U. S.

×

× ×

R V − R B− R

BV RI V − R

B R ∼ k_{R} = 0.11 k_{B} = 0.22 B R R B− R B− R ≃ 1 B− R ≃ 1.5

?? ??

˚ A/pix dispersion) and the 600/7500 grating on the red side (2.33Å/pix dispersion). Our exposure times ranged from 10 min to 50 min for μ_{0} ≃ 20 to μ_{0} ≃ 22 objects, respectively. We performed the standard flatfielding and extraction using the IRAF NOAO LONGSLIT and APEX-TRACT packages. We found that the dichroic produces a lowfrequency ringing in the blueside spectra akin to the fringing patterns requiring removal from the redside. We corrected these effects using local, continuum dome flats taken at the position of every object. To determine the velocity, we crosscorrelated the spectra with a velocitytemplate galaxy observed during the run each night and the IRAF routine FX-COR.

The Keck spectra were taken with LRIS Oke95 in longslit mode, and ESI Sheinis02 in echellette mode. During these runs we focused on the lowest surface brightness candidates, 20μ_{0}24 with exposure times ranging from about 20 to 90 minutes. We used mediumto largesized slits to gather as much light from the LSB candidates as possible. On the red side of LRIS we used the 600/5000, 600/4500, and 600/7500 gratings, and in the most recent run used the new blue side spectrograph, LRIS-B Mc-Carthy98 also Steidel et al. 2003 in prep, with the D560 dichroic and the 600/4000 grism. The LRIS-R spectra were taken in doubleamp readout mode, and flat fielded with continuum lamp spectra where we fit the response function of each amp individually. We fit and removed the slight pincushion distortion using IDEN-TIFY and REI-DEN-TIFY in the IRAF LONGSLIT package to trace out the distortions with arclamps. We determined the 2D transformation of the distortions with TRANS-FORM, and upon applying the transformation we simultaneously determined the wavelength solution. We then used the APEX-TRACT package to extract the calibrated spectrum.

The LRIS-B spectra were taken in fouramp, longslit mode, using the D560 dichroic and the new blue camera. The camera uses two 2kx4k CCDs with a 100 pixel gap between them that corresponds to 135 on the sky. Since the standard longslit position places the object in the detector gap, we used the “slitb” pointing origin which centers the object ∼30 from the gap, on the CCD with higher quantum efficiency. We found the images from the two CCDs were slightly misaligned in the final image, requiring a relative shift of two pixels in the ydirection to align the sky lines across the chip gap. The standard reduction was performed using the IRAF LONGSLIT package, and redshifts were determined using crosscorrelation with FX-COR and a SAO star template unless emission lines were present. For emissionline galaxies, we determined redshifts directly from the emission line identifications.

The ESI spectra from March 2000 were reduced using IRAF routines from the ECHELLE package of tasks, and the April 2003 spectra used the standalone reduction package MAuna Kea Echelle Extraction (MA-

Government grant NAG W-2166. IRAF is distributed by the National Optical Astronomy Observatories, which are operated by the Association of Universities for Research in Astronomy, Inc., under cooperative